# Existence, Sharp Boundary Asymptotics, and Stochastic Optimal Control for Semilinear Elliptic Equations with Gradient-Dependent Terms and Singular Weights

### 08 Mar 2026 v1

@@ 4. Sharp Boundary Asymptotics
The derivation of the limit constants $\xi_{1},\xi_{2}$ relies on a precise asymptotic balance in the equation $\mathcal{L}u=0$ as $d(x)\to 0$. Let $u$ be a solution and assume without loss of generality the gradient-dominant case $q<\beta+2$.
Step 1: Singular order balance. As established in Section 2, for a blow-up rate $\gamma$, the Laplacian $\Delta u$ scales as $d(x)^{-\gamma-2}$, while the weighted gradient term $b(x)|\nabla u|^{q}$ scales as $d(x)^{\beta-q(\gamma+1)}$. Equating the singular orders:
$\gamma-2=\beta-q\gamma-q\implies\gamma(q-1)=\beta-q+2\implies\gamma=\frac{\beta-q+2}{q-1}.$
The reaction term $a(x)u(x)$ behaves like $d(x)^{\alpha-\gamma}$. Since $\alpha>-2$, we have $\alpha-\gamma>-2-\gamma$, which confirms that the reaction term is of lower singular order and vanishes after multiplying the equation by $d(x)^{\gamma+2}$ and taking the limit.
Step 2: Constant identification. Multiply the equation
@@
by $d(x)^{\gamma+2}$. As $x\rightarrow\partial\Omega$, we have $u(x)\sim\xi d(x)^{-\gamma}$. Then:
$-\Delta u$ $\sim-\xi\gamma(\gamma+1)d(x)^{-\gamma-2},$
$b(x)h(|\nabla u|)$ $\sim b(x)l_{0}|\nabla u|^{q}\sim b(x)l_{0}(\xi\gamma)^{q}d(x)^{-q(\gamma+1)}.$
In the limit $d(x)\rightarrow 0$:
$\xi\gamma(\gamma+1)d(x)^{-\gamma-2}+b(x)l_{0}\xi^{q}\gamma^{q}d(x)^{-q(\gamma+1)}\approx 0.$
Using the limits $b_{1},b_{2}$ for $b(x)d(x)^{-\beta}$, we obtain the algebraic equation:
$b_{i}l_{i}\gamma^{q}\xi^{q}-\gamma(\gamma+1)\xi=0.$
Solving for $\xi$ yields the critical values:
$\xi^{q}/\xi=\frac{\gamma(\gamma+1)}{b_{i}l_{i}\gamma^{q}}\implies\xi=\left(\frac{\gamma(\gamma+1)}{b_{i}l_{i}\gamma^{q}}\right)^{\frac{1}{q-1}}.$
The inequalities for $\liminf$ and $\limsup$ follow from the comparison principle applied to the barriers $C_{-}v^{-\gamma}$ and $C_{+}v^{-\gamma}$. If the limits for weights and $h$ exist, the matching of upper and lower barriers forces the existence of the limit $\xi_{0}$. ∎
@@ 5. Stochastic Control and Numerical Verification
We now establish the rigorous connection between the boundary blow-up solution and stochastic optimal control, following the paradigm of Lasry and Lions [17].
5.1. Controlled Diffusion and Admissibility
Let $(\Omega_{0},\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space supporting an $N$-dimensional standard Brownian motion $\{W_{t}\}_{t\geq 0}$. Consider the controlled diffusion
$dX_{t}=\xi_{t}\,dt+\sqrt{2}\,dW_{t},\quad X_{0}=x\in\Omega.$ (5.1)
@@ 5.1. Controlled Diffusion and Admissibility
    The process $X_{t}$ defined by (5.1) remains in $\Omega$ for all time almost surely:
    $\mathbb{P}\left(X_{t}\in\Omega\text{ for all }t\geq 0\right)=1.$
@@
The set $\mathcal{A}_{\mathrm{sc}}(x)$ is nonempty for each $x\in\Omega$. This is a nontrivial fact that follows from the existence of the blow-up solution and the construction of optimal feedback controls below.
@@ 5.2. Legendre Transform and Running Cost
The connection between the PDE (1.1) and stochastic control is mediated by the Legendre transform.
@@
$h^{*}(s)=\sup_{t\geq 0}\{st-h(t)\},\quad s\geq 0.$ (5.2)
By standard convex analysis, $h^{*}$ is strictly convex, $h^{*}(0)=0$, and $h^{*}(s)\asymp s^{q^{\prime}}$ where $q^{\prime}=q/(q-1)$ is the Hölder conjugate exponent.
@@
    Near $\partial\Omega$: $L(x,\xi)\asymp d(x)^{\beta(1-q^{\prime})}|\xi|^{q^{\prime}}$, which blows up as $d(x)\to 0$ for fixed $|\xi|>0$.
@@
Properties (1)–(3) follow from the properties of $h^{*}$. For (4), use $b(x)\asymp d(x)^{\beta}$ and compute
@@
Since $q^{\prime}>1$ and $\beta(1-q^{\prime})<0$ when $\beta>0$ (or appropriately signed), $L$ becomes singular near the boundary. ∎
5.3. The Value Function
Following Lasry and Lions [17], we define the value function for the infinite-horizon state-constrained problem.
@@ 5.3. The Value Function
For $x\in\Omega$, define
$V(x)=\inf_{\xi\in\mathcal{A}_{\mathrm{sc}}(x)}\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a(X_{s})ds}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$ (5.4)
Remark 5.7 (Well-posedness of the integral).
@@
    The discount factor $e^{-\int_{0}^{t}a(X_{s})ds}$ ensures integrability for large $t$ since $a(x)>0$;
@@
    For strictly admissible controls, $X_{t}$ stays away from $\partial\Omega$, so $L(X_{t},\xi_{t})$ remains finite;
@@
    The blow-up of $L$ near $\partial\Omega$ creates an infinite penalty for trajectories approaching the boundary, effectively enforcing the state constraint.
@@ 5.4. Verification Theorem
The unique solution $u\in C^{2}(\Omega)$ from Theorem 1.5 coincides with the value function:
@@
$\xi^{*}(x)=-b(x)(h^{\prime-1}\left(\frac{|\nabla u(x)|}{1}\right)\frac{\nabla u(x)}{|\nabla u(x)|}=-b(x)(h^{*})^{\prime}(|\nabla u(x)|)\frac{\nabla u(x)}{|\nabla u(x)|}.$ (5.6)
@@
Let $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ be any strictly admissible control, and let $X_{t}$ be the corresponding trajectory. For $n\in\mathbb{N}$, define the stopping times
$\tau_{n}=\inf\{t>0:d(X_{t})\leq 1/n\}\wedge n.$
Since $\xi$ is strictly admissible, $\tau_{n}\to\infty$ as $n\to\infty$ almost surely.
Apply Itô’s formula to $e^{-\int_{0}^{t}a(X_{s})ds}u(X_{t})$ on $0,\tau_{n}]$:
$e^{-\int_{0}^{\tau_{n}}a(X_{s})ds}u(X_{\tau_{n}})$
$=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a(X_{s})ds}\left[-a(X_{t})u(X_{t})+\Delta u(X_{t})+\nabla u(X_{t})\cdot\xi_{t}\right]dt$
$\quad+\sqrt{2}\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a(X_{s})ds}\nabla u(X_{t})\cdot dW_{t}.$ (5.7)
Using the PDE
$\Delta u+bh(|\nabla u|)+au=f,$
we substitute:
$\Delta u=bh(|\nabla u|)+au-f.$
Thus (5.7) becomes:
$e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left[bh(|\nabla u|)-f+\nabla u\cdot\xi_{t}\right]dt$
@@
By the Fenchel–Young inequality for convex functions:
@@
$bh(|\nabla u|)+\nabla u\cdot\xi_{t}\geq-L(X_{t},\xi_{t}).$ (5.10)
Taking expectations in (5.8) (the martingale term vanishes):
$\mathbb{E}\left[e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})\right]$ $\geq u(x)-\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$ (5.11)
@@
$u(x)\leq\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left(L+f\right)dt\right]+\mathbb{E}\left[e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})\right].$ (5.12)
As $n\rightarrow\infty$: the first term converges to the full integral by monotone convergence. For the second term, since $\xi$ is strictly admissible, $X_{\tau_{n}}$ stays in $\Omega$, so $u(X_{\tau_{n}})$ is finite. Moreover,
$e^{-\int_{0}^{\tau_{n}}ads}\rightarrow 0\text{ as }\tau_{n}\rightarrow\infty$
(since $a>0$), so this term vanishes.
@@
$u(x)\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}ads}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$
@@
Consider the feedback control defined by (5.6). At this choice, the Fenchel inequality (5.9) becomes an equality:
@@
This is the first-order optimality condition for the Legendre transform.
Let $X_{t}^{*}$ be the trajectory under the feedback control $\xi_{t}^{*}=\xi^{*}(X_{t}^{*})$. The crucial observation is that the blow-up behavior of $u$ near $\partial\Omega$ implies that $|\xi^{*}(x)|\to\infty$ as $x\to\partial\Omega$. Specifically, since $\nabla u$ blows up as $d(x)^{-\gamma-1}$, the drift $\xi^{*}$ points strongly into the interior of $\Omega$ near the boundary.
@@
$\mathbb{P}\left(X_{t}^{*}\in\Omega\text{ for all }t\geq 0\right)=1.$
@@
Repeating the Itô calculation with equality in (5.9):
$u(x)=\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}ads}\left(L(X_{t}^{*},\xi_{t}^{*})+f(X_{t}^{*})\right)dt\right]\geq V(x).$
@@
Theorem 5.8 is a generalization of Theorem 1.1 in [17] to equations with weight functions $a(x)$ and $b(x)$. The original Lasry–Lions result treats the case $a=\lambda\in\left(0,\infty\right)$, $b=1$, and $h(|\nabla u(x)|)=|\nabla u(x)|^{p}$ for $p>1$. Our formulation incorporates:
@@
    A discount rate $a(x)$ in the cost functional;

### 22 Mar 2026 v2

@@ 5. Sharp Boundary Asymptotics
The derivation of the limit constants $\xi_{1},\xi_{2}$ relies on a precise asymptotic balance in the equation $\mathcal{L}u=0$ as $d(x)\to 0$. Let $u$ be the unique solution from Theorem 2.9 and assume the gradient-dominant case $q<\beta+2$.
Step 1: Singular order balance.
As established in Section 3, for a blow-up rate $\gamma$, the individual terms in the equation scale as follows near $\partial\Omega$:
$\Delta u$ $\sim C\,\gamma\,(\gamma+1)\,d(x)^{-\gamma-2},$
$b(x)\,h(|\nabla u|)$ $\sim b(x)\,l_{0}\,(C\,\gamma)^{q}\,d(x)^{-q(\gamma+1)}$
$\sim b_{0}\,l_{0}\,(C\,\gamma)^{q}\,d(x)^{\beta-q(\gamma+1)},$
$a(x)\,u$ $\sim a_{0}\,C\,d(x)^{\alpha-\gamma},$
$f(x)$ $=O(1).$
Equating the singular orders of the two leading terms (Laplacian and weighted gradient):
$\gamma-2=\beta-q\gamma-q\implies\gamma(q-1)=\beta-q+2\implies\gamma=\frac{\beta-q+2}{q-1}.$ (5.1)
The reaction term a(x)u(x)a(x)\,u(x) behaves like $d(x)^{\alpha-\gamma}$. Since $\alpha>-2$ by Assumption 2.6, we have $\alpha-\gamma>-2-\gamma=-(\gamma+2)$. This confirms that the reaction term is of lower singular order than the leading terms $d(x)^{-\gamma-2}$ and vanishes after multiplying the equation by $d(x)^{\gamma+2}$ and taking the limit $d\to 0$.
Step 2: Constant identification.
Multiply the equation
@@
by $d(x)^{\gamma+2}$. As $x\to\partial\Omega$, let $\xi=\lim_{d\to 0}u(x)\,d(x)^{\gamma}$ (assuming the limit exists). Then $u(x)\sim\xi\,d(x)^{-\gamma}$, and:
$-\Delta u$ $\sim-\xi\,\gamma\,(\gamma+1)\,d(x)^{-\gamma-2},$ (5.2)
$b(x)\,h(|\nabla u|)$ $\sim b(x)\,l_{0}\,(\xi\,\gamma)^{q}\,d(x)^{-q(\gamma+1)}.$ (5.3)
After multiplying by $d^{\gamma+2}$ and using (5.1), the equation becomes in the limit:
$\xi\,\gamma\,(\gamma+1)+b_{i}\,l_{i}\,(\xi\,\gamma)^{q}=0,$ (5.4)
where the subscript $i$ indicates whether we use the liminf or limsup of $b(x)\,d(x)^{-\beta}$ and $h(s)\,s^{-q}$.
Solving (5.4) for $\xi$:
$b_{i}\,l_{i}\,\gamma^{q}\,\xi^{q}=\gamma\,(\gamma+1)\,\xi\implies\xi^{q-1}=\frac{\gamma\,(\gamma+1)}{b_{i}\,l_{i}\,\gamma^{q}}\implies\xi=\left(\frac{\gamma\,(\gamma+1)}{b_{i}\,l_{i}\,\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (5.5)
Step 3: Liminf–limsup estimates.
The inequalities for $\liminf$ and $\limsup$ in (2.6) follow from the comparison principle applied to the barriers $C_{-}\,v^{-\gamma}$ and $C_{+}\,v^{-\gamma}$ from Lemma 3.2. Since $C_{-}\leq u\,d^{\gamma}\leq C_{+}$ and the constants $C_{\pm}$ can be made arbitrarily close to $C^{*}$ (from above and below), we obtain the precise bounds $\xi_{1}$ and $\xi_{2}$ by using the extremal values $b_{2},l_{2}$ (for $\xi_{1}$, the lower bound) and $b_{1},l_{1}$ (for $\xi_{2}$, the upper bound).
If both limits $b_{0}=\lim_{d\to 0}b(x)\,d^{-\beta}$ and $l_{0}=\lim_{s\to\infty}h(s)\,s^{-q}$ exist, then $b_{1}=b_{2}=b_{0}$ and $l_{1}=l_{2}=l_{0}$, which forces $\xi_{1}=\xi_{2}=\xi_{0}$, establishing (2.8). ∎
Remark 5.1 (The Universal Correction Constant).
The constant $\xi_{0}$ defined in (2.8) has a direct economic interpretation. In the stochastic control framework (Section 6), $\xi_{0}$ determines the intensity of the “boundary penalty” that the optimal strategy must impose to keep the system within $\Omega$. A larger $\xi_{0}$ (corresponding to smaller $b_{0}$ or $l_{0}$, i.e., weaker singular weights or weaker gradient cost) means the controller must exert a stronger corrective effort near the boundary. Conversely, a smaller $\xi_{0}$ (stronger singular weights) means the natural cost structure already provides significant deterrence, reducing the required corrective effort.
@@ 6. Stochastic Control and Numerical Verification
In this section, we develop the complete stochastic control framework that gives rise to the mathematical problem studied. We provide all details leading from the controlled diffusion model to the Hamilton–Jacobi–Bellman equation (1.3), and we prove the verification theorem establishing the equivalence between the PDE solution and the value function.
6.1. The Controlled Diffusion Model
Let $(\Omega_{0},\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space supporting an $N$-dimensional standard Brownian motion $\{W_{t}\}_{t\geq 0}$, satisfying the usual conditions (completeness and right-continuity of the filtration).
Consider a system whose state $X_{t}\in\mathbb{R}^{N}$ evolves according to the controlled stochastic differential equation (SDE):
$dX_{t}=\xi_{t}\,dt+\sqrt{2}\,dW_{t},\quad X_{0}=x\in\Omega,$ (6.1)
where $\xi_{t}$ is the control process (representing the corrective action taken by the decision-maker) and the factor $\sqrt{2}$ is chosen so that the generator of the uncontrolled diffusion is exactly the Laplacian $\Delta$ (rather than $\frac{1}{2}\Delta$).
The derivation of the generator proceeds as follows. For a smooth function $\phi\in C^{2}(\Omega)$, Itô’s formula gives:
$d\phi(X_{t})$ $=\nabla\phi(X_{t})\cdot dX_{t}+\frac{1}{2}\,\mathrm{tr}\bigl(D^{2}\phi(X_{t})\,(\sqrt{2}\,I)(\sqrt{2}\,I)^{T}\bigr)\,dt$
$=\bigl[\nabla\phi(X_{t})\cdot\xi_{t}+\Delta\phi(X_{t})\bigr]\,dt+\sqrt{2}\,\nabla\phi(X_{t})\cdot dW_{t}.$ (6.2)
The infinitesimal generator of the uncontrolled process ($\xi\equiv 0$) is therefore $\mathcal{A}\phi=\Delta\phi$, as claimed.
6.2. Admissibility and State Constraints
The fundamental constraint is that the state must remain in $\Omega$ at all times.
@@ 6.2. Admissibility and State Constraints
    The SDE (6.1) has a unique strong solution;
  3. (3)
    The process $X_{t}$ remains in $\Omega$ for all time almost surely:
    $\mathbb{P}\bigl(X_{t}\in\Omega\text{ for all }t\geq 0\bigr)=1.$ (6.3)
@@
The set $\mathcal{A}_{\mathrm{sc}}(x)$ is nonempty for each $x\in\Omega$. This is a nontrivial fact that follows from the existence of the blow-up solution and the construction of optimal feedback controls below (Theorem 6.9).
@@ 6.3. Legendre Transform and Running Cost
The connection between the PDE (1.3) and stochastic control is mediated by the Legendre–Fenchel transform (convex conjugation).
@@
$h^{*}(s)=\sup_{t\geq 0}\bigl\{s\,t-h(t)\bigr\},\quad s\geq 0.$ (6.4)
By standard convex analysis (see, e.g., Rockafellar [25]), $h^{*}$ has the following properties:
Lemma 6.4 (Properties of the conjugate).
Under Assumption 2.4:
  1. (1)
    $h^{*}$ is strictly convex, $h^{*}(0)=0$, and $h^{*}\in C^{1}(0,\infty)$;
  2. (2)
    $h^{*}(s)\asymp s^{q^{\prime}}$ where $q^{\prime}=q/(q-1)$ is the Hölder conjugate exponent;
  3. (3)
    The supremum in (6.4) is achieved at $t=(h^{\prime})^{-1}(s)$, and $(h^{*})^{\prime}(s)=(h^{\prime})^{-1}(s)$;
  4. (4)
    The Fenchel–Young inequality holds: $s\,t\leq h(t)+h^{*}(s)$ for all $s,t\geq 0$, with equality if and only if $s=h^{\prime}(t)$.
@@
    Near $\partial\Omega$: $L(x,\xi)\asymp d(x)^{\beta(1-q^{\prime})}\,|\xi|^{q^{\prime}}$, which blows up as $d(x)\to 0$ for fixed $|\xi|>0$ provided $\beta>0$.
@@
Properties (1)–(3) follow directly from the properties of $h^{*}$ (Lemma 6.4) and the positive homogeneity of the scaling $|\xi|/b(x)$. For property (4), we use $b(x)\asymp d(x)^{\beta}$ from Assumption 2.6 (W2) and compute:
@@
Since $q^{\prime}>1$, we have $1-q^{\prime}<0$, so $\beta(1-q^{\prime})<0$ when $\beta>0$, ensuring $L(x,\xi)\to\infty$ as $d(x)\to 0$.
The crucial duality identity is: for each $x\in\Omega$,
$b(x)\,h(|p|)=\sup_{\xi\in\mathbb{R}^{N}}\bigl\{-\xi\cdot p-L(x,\xi)\bigr\}.$ (6.6)
This follows from the definition of $h^{*}$ and the substitution $\xi=-b(x)\,h^{\prime}(|p|)\,p/|p|$. ∎
6.4. Derivation of the Hamilton–Jacobi–Bellman Equation
We now derive the HJB equation (1.3) from the stochastic control problem via the dynamic programming principle. This derivation makes explicit how each term in the PDE arises from the control problem.
@@ 6.4. Derivation of the Hamilton–Jacobi–Bellman Equation
For $x\in\Omega$, define the value function
$V(x)=\inf_{\xi\in\mathcal{A}_{\mathrm{sc}}(x)}\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a(X_{s})\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$ (6.7)
Remark 6.8 (Well-posedness of the cost functional).
@@
    The discount factor $e^{-\int_{0}^{t}a(X_{s})\,ds}$ ensures integrability for large $t$ since $a(x)>0$ everywhere in $\Omega$ (Assumption 2.6).
@@
    For strictly admissible controls, $X_{t}$ stays in $\Omega$, so $L(X_{t},\xi_{t})$ remains finite for each $t$.
@@
    The blow-up of $L$ near $\partial\Omega$ (Lemma 6.6, property (4)) creates an infinite penalty for trajectories approaching the boundary, effectively enforcing the state constraint.
Derivation via the Dynamic Programming Principle. Assume $V$ is smooth. For a small time increment $\Delta t>0$, the dynamic programming principle gives:
$V(x)=\inf_{\xi}\,\mathbb{E}\left[\int_{0}^{\Delta t}e^{-\int_{0}^{t}a(X_{s})\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt+e^{-\int_{0}^{\Delta t}a(X_{s})\,ds}\,V(X_{\Delta t})\right].$ (6.8)
Expanding $V(X_{\Delta t})$ via Itô’s formula (6.2) and taking $\Delta t\to 0$, we obtain:
$0=\inf_{\xi\in\mathbb{R}^{N}}\left\{\Delta V(x)+\nabla V(x)\cdot\xi-a(x)\,V(x)+L(x,\xi)+f(x)\right\}.$ (6.9)
The infimum over $\xi$ can be computed explicitly using (6.6). Using the identity $\inf_{\xi}\{\nabla V\cdot\xi+L(x,\xi)\}=-\sup_{\xi}\{-\nabla V\cdot\xi-L(x,\xi)\}=-b(x)\,h(|\nabla V|)$, we obtain:
$0=\Delta V-b(x)\,h(|\nabla V|)-a(x)\,V+f(x),$ (6.10)
which is precisely equation (1.3) with $u=V$.
This derivation shows explicitly how each term in (1.3) arises:
  • $\Delta V$: the generator of the Brownian noise $\sqrt{2}\,dW_{t}$;
  • $b(x)\,h(|\nabla V|)$: the optimal cost of corrective action, via Legendre duality;
  • a(x)Va(x)\,V: the state-dependent discounting;
  • $f(x)$: the external source/cost.
@@ 6.5. Verification Theorem
Under Assumptions 2.12.6, the unique solution $u\in C^{2}(\Omega)$ from Theorem 2.9 coincides with the value function:
@@
$\xi^{*}(x)=-b(x)\,h^{\prime}(|\nabla u(x)|)\,\frac{\nabla u(x)}{|\nabla u(x)|}=-b(x)\,(h^{*})^{\prime}(|\nabla u(x)|)^{-1}\,\frac{\nabla u(x)}{|\nabla u(x)|}.$ (6.12)
The identity $(h^{\prime})^{-1}=(h^{*})^{\prime}$ is a standard consequence of the Legendre transform: the supremum in $h^{*}(s)=\sup_{t}\{st-h(t)\}$ is achieved at $s=h^{\prime}(t)$, so $t=(h^{\prime})^{-1}(s)$. Differentiating the identity $h^{*}(s)=s\,(h^{\prime})^{-1}(s)-h((h^{\prime})^{-1}(s))$ with respect to $s$ yields $(h^{*})^{\prime}(s)=(h^{\prime})^{-1}(s)$.
@@
Let $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ be any strictly admissible control, and let $X_{t}$ be the corresponding trajectory satisfying (6.1). For $n\in\mathbb{N}$, define the stopping times
$\tau_{n}=\inf\{t>0:d(X_{t})\leq 1/n\}\wedge n.$ (6.13)
Since $\xi$ is strictly admissible, $X_{t}$ remains in $\Omega$ almost surely for all $t\geq 0$. For any $T>0$, the path $t\mapsto X_{t}$ is continuous and the image $X([0,T])$ is a compact subset of $\Omega$, hence $\inf_{t\in[0,T]}d(X_{t})>0$ almost surely. This ensures that for almost all $\omega$, there exists $n_{0}(\omega)$ such that $\tau_{n}(\omega)=n$ for all $n\geq n_{0}(\omega)$, giving $\tau_{n}\to\infty$ as $n\to\infty$ almost surely.
Apply Itô’s formula to $e^{-\int_{0}^{t}a(X_{s})\,ds}\,u(X_{t})$ on $0,\tau_{n}]$. By (6.2):
$e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})$
$=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl[-a(X_{t})\,u(X_{t})+\Delta u(X_{t})+\nabla u(X_{t})\cdot\xi_{t}\bigr]\,dt$
$\quad+\sqrt{2}\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\nabla u(X_{t})\cdot dW_{t}.$ (6.14)
Using the PDE $\Delta u+b\,h(|\nabla u|)+a\,u=f$, we substitute $\Delta u=b\,h(|\nabla u|)+a\,u-f$. Thus (6.14) becomes:
$e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})$ $=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl[b\,h(|\nabla u|)-f+\nabla u\cdot\xi_{t}\bigr]\,dt$
@@
By the Fenchel–Young inequality (Lemma 6.4, property (4)):
@@
$b\,h(|\nabla u|)+\nabla u\cdot\xi_{t}\geq-L(X_{t},\xi_{t}).$
Taking expectations in (6.15) (the Itô integral is a martingale with zero expectation, since $u$ and $\nabla u$ are bounded on the compact set $\{x:d(x)\geq 1/n\}$):
$\mathbb{E}\bigl[e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})\bigr]\geq u(x)-\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$ (6.17)
@@
$u(x)\leq\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl(L+f\bigr)\,dt\right]+\mathbb{E}\bigl[e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})\bigr].$ (6.18)
As $n\to\infty$: the first term converges to the full integral by monotone convergence. For the second term, since $\xi$ is strictly admissible, $X_{\tau_{n}}$ stays in $\Omega$, so $u(X_{\tau_{n}})$ is finite. Moreover, $e^{-\int_{0}^{\tau_{n}}a\,ds}\to 0$ as $\tau_{n}\to\infty$ (since $a>0$), so this term vanishes.
@@
$u(x)\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$
@@
Consider the feedback control defined by (6.12). At this choice, the Fenchel–Young inequality (6.16) becomes an equality:
@@
This is the first-order optimality condition for the Legendre transform: the supremum in $h^{*}$ is achieved.
Let $X_{t}^{*}$ be the trajectory under the feedback control $\xi_{t}^{*}=\xi^{*}(X_{t}^{*})$. The blow-up behavior of $u$ near $\partial\Omega$ implies $|\xi^{*}(x)|\to\infty$ as $x\to\partial\Omega$. Specifically, since $\nabla u$ blows up as $d(x)^{-\gamma-1}$, the drift $\xi^{*}$ points strongly into the interior of $\Omega$ near the boundary.
@@
$\mathbb{P}\bigl(X_{t}^{*}\in\Omega\text{ for all }t\geq 0\bigr)=1.$
@@
Repeating the Itô calculation with equality in (6.16) (i.e., using (6.19)):
$u(x)=\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t}^{*},\xi_{t}^{*})+f(X_{t}^{*})\bigr)\,dt\right]\geq V(x).$
@@
Theorem 6.9 generalizes Theorem 1.1 in [17] to equations with state-dependent weight functions $a(x)$ and $b(x)$. The original Lasry–Lions result treats the case $a=\lambda\in(0,\infty)$ (constant discount rate), $b=1$ (no spatial weight), and $h(|\nabla u|)=|\nabla u|^{p}$ for $p>1$. Our formulation incorporates:
@@
    A state-dependent discount rate $a(x)$ in the cost functional;
@@
    A general strictly convex Hamiltonian $h$ with power-type growth;
  4. (4)